# She’s Got Her Mother’s Hair: End-to-End Collapsar Simulations Unveil the Origin of Black Holes’ Magnetic Field

### v1

@@ Abstract
Relativistic jets from a black hole (BH) following the core-collapse of a massive star (“collapsar”) is a leading model for gamma-ray bursts (GRBs). However, the two key ingredients for a Blandford-Znajek (BZ) powered jet -- rapid rotation and a strong magnetic field -- seem mutually exclusive. Strong fields in the progenitor star’s core transport angular momentum outwards more quickly, slowing down the core before collapse. Using MESA stellar evolution models followed to core-collapse, we explicitly show that the small length-scale of the instabilities, likely responsible for angular momentum transport in the core (e.g., Tayler-Spruit), results in a low net magnetic flux fed to the BH horizon, far too small to power GRB jets. Instead, we propose a novel scenario in which collapsar BHs acquire their magnetic “hair” from their progenitor proto-neutron star (PNS), which is likely highly magnetized from an internal dynamo. We evaluate the conditions for the BH accretion disk to pin the PNS magnetosphere to its horizon immediately after the collapse. Our results show that the PNS’s pre-collapse energy matches the excess energy in Ic-BL supernovae, while the nascent BH’s spin and magnetic flux produce jets consistent with observed GRB characteristics. We map our MESA models to 3D general-relativistic magnetohydrodynamic simulations and confirm that accretion disks confine the strong magnetic flux initiated near a rotating BH, enabling the launch of successful GRB jets, whereas a slower spinning BH or one without a disk fails to do so. Our model indicates that standard GRB jets emerge whenever an accretion disk forms.
@@ 1 Introduction
Massive star collapse constitutes the primary formation channel for neutron stars (NSs) and black holes (BHs) in the Universe (e.g., Janka 2012; Burrows & Vartanyan 2021; Soker 2024). Observations of distant supernovae (SNe) and their remnants in our Galaxy indicate that a significant fraction of NSs are born as “magnetars”, with dipole magnetic field strengths of $B\approx 10^{14}-10^{15}\,\,{{\rm G}}$ (Kaspi & Beloborodov, 2017; Beniamini et al., 2019), some of which may be rapidly spinning millisecond (ms) magnetars (Kasen & Bildsten, 2010; Mösta et al., 2015). Based on the light curves of stripped-envelope SN (SESN), it has been argued that all SESNe, and possibly all SNe, give birth to either a ms-magnetar or an accreting BH (Woosley et al., 2021; Rodríguez et al., 2024).
In a small fraction of collapsing stars, the newly formed compact object powers a magnetized relativistic jet capable of generating a gamma-ray burst (GRB). Whether the electromagnetically-driven jets are launched by a BH (Blandford & Znajek 1977, hereafter BZ; Popham et al. 1999; MacFadyen & Woosley 1999) or a NS (Goldreich & Julian, 1969; Usov, 1992; Thompson, 1994; Metzger et al., 2007, 2011), they necessitate rapid rotation and an extremely strong large-scale poloidal magnetic field $\gtrsim 10^{15}$ G to facilitate jet launching (e.g., Tchekhovskoy & Giannios 2015; Gottlieb et al. 2022a). This raises the question of whether a rapidly rotating star, whose core prior to collapse is predicted to possess a comparatively weak magnetic field $B\approx 10^{10}\,\,{{\rm G}}$ (Aguilera-Dena et al., 2018; Müller, 2024), can generate such strongly magnetized compact objects.
Two possible magnetar formation channels have been proposed (e.g., Spruit 2008; Igoshev et al. 2021 for reviews). In the “fossil field” scenario, a NS acquires its magnetic field externally, from the collapsing star (Ferrario & Wickramasinghe, 2006). If the field has a substantial poloidal component, magnetic flux freezing during the collapse will amplify the field by a factor of $(R_{c}/\,{{R_{\rm NS}}})^{2}$, where $R_{c}$ is the radius of the precollapse star enclosing the NS mass, and $\,{{R_{\rm NS}}}$ is the NS radius. This factor can be $\gtrsim 10^{5}$, sufficient to generate a NS magnetic field $\gtrsim 10^{15}\,\,{{\rm G}}$ (see e.g., Makarenko et al., 2021). However, the topology of the core’s magnetic field as predicted by, e.g., the Tayler-Spruit dynamo (TSD) (Tayler, 1973; Spruit, 2002; Fuller et al., 2019; Skoutnev & Beloborodov, 2024) is primarily toroidal, such that only a small fraction of the total progenitor flux will contribute to the dipole field of the newly formed proto-neutron star (PNS). Furthermore, even if the poloidal field is relatively strong, the small spatial scales that characterize the TSD render the field unlikely to possess a uniform polarity prior to or after the collapse, resulting in a reduced net flux reaching small scales.
@@
At least three mechanisms can be at play over the first few seconds of the PNS cooling evolution: (1) the magnetorotational instability (MRI; Balbus & Hawley, 1991), particularly in the outer layers of the PNS where the angular velocity decreases outward (Akiyama et al., 2003; Obergaulinger et al., 2009; Reboul-Salze et al., 2021); (2) the TSD, particularly in stably stratified regions close to the PNS core where the MRI cannot operate (e.g., Margalit et al. 2022); and (3) an $\alpha-\omega$ dynamo in the convective regions of the PNS (e.g., Duncan & Thompson, 1992; Thompson & Duncan, 1993; Bonanno et al., 2003; Raynaud et al., 2020; White et al., 2022; Barrère et al., 2024). All three mechanisms are in principle capable of generating fields of strength $\lesssim 10^{16}\,\,{{\rm G}}$ (Thompson & Murray, 2001; Akiyama et al., 2003), though only the MRI and a convective dynamo are likely to dominate near the surface of the PNS (however, see Barrère et al. 2022). The Rossby number $Ro=P/\tau_{c}$ determines the relative contribution of these mechanisms, where $P$ is the PNS spin period and $\tau_{c}$ is the convective turnover time in its envelope. When rotation is slower than the convective motions such that $Ro\gtrsim 1$, the MRI dominates as it grows on the rotation timescale, leading to a stochastic dynamo which generates small-scale multipolar fields (e.g., Guilet et al. 2022; Jacquemin-Ide et al. 2023). On the other hand, for sufficiently rapid rotation $P\lesssim 10$ ms and $Ro\lesssim 0.2$, the MRI is subdominant to the $\alpha-\omega$ dynamo, which can generate a strong ordered dipolar field (Nagakura et al., 2020; Raynaud et al., 2020; White et al., 2022), whose magnitude grows with the rotation rate (Masada et al., 2022).
@@
While the origin of the NS magnetic fields has been extensively debated in the literature, the question of how BHs acquire enough magnetic flux to power relativistic jets in jetted core-collapse events like GRBs has received less attention. Unlike NSs, BHs cannot generate their own fields, which must be inherited externally. One possibility is that the BH acquires its magnetic flux directly through accretion from the infalling progenitor envelope (e.g., Tchekhovskoy & Giannios 2015). However, as we will show, the same arguments raised above against magnetars acquiring large-scale fields from the progenitor star – namely the strength and small coherence scale of the precollapse poloidal field – also apply to the accreting BH case, particularly for stellar cores that retain enough angular momentum to produce GRB jets in the first place. Another possibility is an amplification of the field through dynamo processes in the accretion disk (e.g., Jacquemin-Ide et al., 2023; Most, 2023). This mechanism might also be inefficient for collapsar GRBs given the expected properties of collapsar disks, as we discuss later in light of our stellar evolution models. On the other hand, most stellar-mass BHs were previously PNS, prior to accreting enough mass to undergo gravitational collapse (e.g., Dessart et al. 2012; Aloy & Obergaulinger 2021). This motivates the possibility that collapsar BHs inherit their magnetic fields from their progenitor PNS.
@@
However, this “balding” process may be postponed if the inherited magnetic field can be pinned to the BH by an external current source, such as that supplied by an accretion disk. While not expected to form around an old NS that collapses in isolation (e.g., accretion-induced collapse in a compact binary; Margalit et al. 2015), such a disk may be present around a differentially-rotating PNS embedded within the infalling envelope of a rotating star. Nevertheless, it remains unaddressed whether, in astrophysical systems such as collapsars or post-binary NS mergers, the BH can maintain the inherited magnetic flux long enough to launch a sustained relativistic jet necessary to power GRBs.
@@
Here, we show explicitly that massive stars with rapidly rotating cores at the time of collapse generate an accretion disk around the central compact object but do not provide sufficient magnetic flux to power a GRB-like jet. Instead, we propose that such events first create a rapidly spinning PNS surrounded by an accretion disk and that internal dynamo processes amplify the PNS magnetic field to magnetar-level strengths. Then, when the magnetar accretes sufficient mass to collapse, its strong poloidal component is retained by the resulting BH due to confinement by the accretion disk. Therefore, even without ongoing accreted magnetic flux, the BH retains its dipolar field which can launch the BZ jets that generate GRBs. The astrophysical setup and sequence of events are illustrated in Figure 1.
@@
The paper is organized as follows. In §2, we begin by presenting stellar evolution models that lead to the formation of a PNS with an accretion disk and use them to explain why the intrinsic poloidal flux fed by stellar infall is likely insufficient to power a GRB jet. In §3, we estimate the conditions necessary for the BH to capture the field of the collapsing PNS, which constrain the expected natal spin and magnetic flux acquired by the BH during this process. In §4, we conduct general-relativistic magnetohydrodynamic (GRMHD) simulations, demonstrating that BHs retain their magnetic field without further flux accumulation, so long as an accretion disk confines them. We conclude in §5, and discuss the implications of our results on GRBs and their associated SNe in §6.
@@
Figure 1: Not-to-scale illustration of the proposed model. Top: A rotating massive star (light blue) collapses, forming a rapidly rotating proto-neutron star (white) surrounded by an accretion disk (dark blue) due to the excess angular momentum of the innermost layers of the stellar core (§2). The magnetic field present in the progenitor star due to the Tayler-Spruit dynamo is too weak and/or small-scale (randomly oriented cyan loops) to contribute a substantial magnetic flux onto the compact object (§2.1). A convective dynamo within the rapidly spinning proto-neutron star efficiently generates a strong, ordered magnetic field (pink lines). Bottom: The proto-neutron star collapses into a moderately spinning black hole (black). The accretion disk surrounding the proto-neutron star at the time of collapse quickly spreads inwards viscously to the black hole’s innermost stable circular orbit, anchoring the proto-neutron star’s magnetic field lines to the BH faster than the “balding” process. This enables the black hole to form with strong magnetic fields already threading through it (§3). The strongly magnetized and moderately spinning black hole then launches a pair of Blandford-Znajek jets (yellow) that break through the outer layers of the collapsing star and powers gamma-ray burst emission (§4).
@@ 2 Black hole formation in stellar evolution models
Figure 2: From top to bottom, as a function of the local free fall timescale: specific angular momentum profile (gray band brackets $j_{\mathrm{ISCO}}(a_{\mathrm{BH}}=0)$ and $j_{\mathrm{ISCO}}(a_{\mathrm{BH}}=1)$); mass density (red) and enclosed mass (orange, right vertical axis); estimate of the accretion rate (magenta) and local infall velocity profile (cyan, right vertical axis); B-field components predicted by the Tayler-Spruit dynamo; Alfvén, TS and free fall local timescales (purple, yellow, and dashed gray, respectively); and estimated cumulative magnetic flux $\Phi$, where the red band is the required flux to power a GRB jet (see text).
@@
The top panel of Fig. 2 depicts the specific angular momentum of the infalling matter, $j=r\omega^{2}$ (blue), and the innermost stable circular orbit (ISCO) specific angular momentum, $j_{\mathrm{ISCO}}$ (black), as a function of the enclosed mass and angular momentum calculated following Bardeen et al. (1972). In the innermost layers, where $\tau_{\mathrm{free\ fall}}\lesssim 3\,\mathrm{s}$, we find $j\gtrsim j_{\mathrm{ISCO}}$, which prohibits a prompt collapse to a BH, and necessitates a phase of a PNS with an accretion disk (see e.g., Dessart et al., 2008; Rahman et al., 2022). This result is sensitive to uncertain angular momentum transport processes in the star, which are virtually unconstrained in this mass regime. As outlined in Appendix A, we adopt angular momentum transport via TSD in radiative layers (Spruit, 2002), which provides relatively weak transport by magnetic torques (compared to Fuller et al. 2019; Fuller & Lu 2022). In convective layers angular momentum transport is assumed to be quite efficient, with a diffusion coefficient proportional to the convective diffusivity. We also neglect possible late spin-up of the core due to gravity waves excited by convective shell burning (e.g. Fuller et al., 2015).
@@
Using the same initial conditions and incorporating magnetic torques from [^1] Implemented in https://github.com/MESAHub/mesa-contrib/. Fuller & Lu (2022), the model still exhibits chemically homogeneous evolution. However, most of the angular momentum is lost post-hydrogen core burning and no layer retains enough angular momentum at the onset of collapse to form an accretion disk. Different assumptions can result in the inner regions spinning less rapidly, allowing a prompt collapse to a BH (MacFadyen & Woosley, 1999). In such cases, one can estimate the BH seed formation time $t_{\mathrm{BH}}$ as the free fall time of the layers with $j<j_{\mathrm{ISCO}}(a(m,j))$, and the accretion disk might form from the outer parts of the stellar envelope. This raises difficulties for this scenario to produce GRB progenitors.
@@
The second panel portrays the mass density profile (red, left vertical axis), and the enclosed mass (orange, right vertical axis). The total mass of the star at the onset of core collapse is $\sim{}25\,M_{\odot}$ due to wind mass loss (see Appendix A). The third panel shows the infall velocity profile (cyan, right axis) – we evolve the star until it reaches an infall velocity of $-300\,\mathrm{km\ s^{-1}}$. The third panel also shows a rough estimate of the mass accretion rate calculated as $\dot{M}={\rm d}m/{\rm d}\tau_{\mathrm{free\ fall}}$ (magenta). The horizontal blue shaded band highlights $0.1\leq\dot{M}/M_{\odot}\,\,{{\rm s}}^{-1}\leq 3$ (see §3).
@@
The fourth panel depicts the radial ($B_{r}$, pink) and azimuthal ($B_{\varphi}$, green) components of the magnetic field $B$ predicted by the TSD (Spruit, 2002). The region where the field drops below 100 G is convective at the onset of core collapse and thus does not contain TSD-generated fields. These fields are, by definition, local, and do not maintain the large-scale structure or a poloidal topology, both of which are required to power electromagnetically-driven jets (see §2.1).
@@
$\tau_{\mathrm{TS}}=\tau_{\mathrm{Alfven}}\times\frac{\omega}{N}\,,$ (3)
where $\omega$ is the local rotational frequency and $N$ is the Brunt-Vaisala frequency (Väisälä, 1925). Since $\omega\lesssim N$ throughout, $\tau_{\mathrm{TS}}\lesssim\tau_{\mathrm{Alfven}}$. In the innermost layers, with $m\lesssim M_{\odot}$ and $\tau_{\mathrm{free\ fall}}\lesssim 0.1\,\mathrm{s}$, both timescales are significantly longer than the free-fall timescale, justifying the assumption of flux-freezing. Moving outward in the region $\,{M_{\odot}}\lesssim m\lesssim 3\,\,{M_{\odot}}$, corresponding to infall times $\tau_{\mathrm{free\ fall}}\lesssim 2\,\mathrm{s}$, the Alfvén and TSD timescales become shorter than the infall timescale. Thus, magnetic diffusion and reconnection could occur, leading to an even lower magnetic flux reaching the seed BH (in the absence of dynamos). However, with $\tau_{\rm TS}\lesssim\tau_{\mathrm{Alfven}}$, the TSD may also regenerate new local B-fields during the collapse, but those would still be local and not have a large-scale poloidal topology.

### v2

# She’s Got Her Mother’s Hair: Unveiling the Origin of Black Hole Magnetic Fields through Stellar to Collapsar Simulations

@@ Abstract
Relativistic jets from a Kerr black hole (BH) following the core collapse of a massive star (“collapsar”) is a leading model for gamma-ray bursts (GRBs). However, the two key ingredients for a Blandford-Znajek powered jet -- rapid rotation and a strong magnetic field -- seem mutually exclusive. Strong fields in the progenitor star’s core transport angular momentum outwards more quickly, slowing down the core before collapse. Through innovative multidisciplinary modeling, we first use MESA stellar evolution models followed to core collapse, to explicitly show that the small length-scale of the instabilities – likely responsible for angular momentum transport in the core (e.g., Tayler-Spruit) – results in a low net magnetic flux fed to the BH horizon, far too small to power GRB jets. Instead, we propose a novel scenario in which collapsar BHs acquire their magnetic “hair” from their progenitor proto-neutron star (PNS), which is likely highly magnetized from an internal dynamo. We evaluate the conditions for the BH accretion disk to pin the PNS magnetosphere to its horizon immediately after the collapse. Our results show that the PNS spin-down energy released before collapse matches the kinetic energy of Type Ic-BL supernovae, while the nascent BH’s spin and magnetic flux produce jets consistent with observed GRB characteristics. We map our MESA models to 3D general-relativistic magnetohydrodynamic simulations and confirm that accretion disks confine the strong magnetic flux initiated near a rotating BH, enabling the launch of successful GRB jets, whereas a slower-spinning BH or one without a disk fails to do so.
@@ 1 Introduction
Massive star collapse constitutes the primary formation channel for neutron stars (NSs) and black holes (BHs) in the Universe (e.g., Janka 2012; Burrows & Vartanyan 2021; Soker 2024). Observations of distant supernovae (SNe) and their remnants in our Galaxy indicate that a significant fraction of NSs are born as “magnetars,” with dipole magnetic field strengths of $B\approx 10^{14}-10^{15}\,\,{{\rm G}}$ (Kaspi & Beloborodov, 2017; Beniamini et al., 2019), some of which may be rapidly spinning millisecond (ms) magnetars (Kasen & Bildsten, 2010; Mösta et al., 2015). Based on the light curves of stripped-envelope SNe (SESNe), it has been argued that all SESNe, and possibly all SNe, give birth to either a millisecond (ms) magnetar or an accreting BH (Woosley et al., 2021; Rodríguez et al., 2024).
In a small fraction of collapsing stars, the newly formed compact object powers a magnetized relativistic jet capable of generating a gamma-ray burst (GRB). Whether the electromagnetically-driven jets are launched by a BH (Blandford & Znajek 1977, hereafter BZ; Popham et al. 1999; MacFadyen & Woosley 1999) or an NS (Goldreich & Julian, 1969; Usov, 1992; Thompson, 1994; Metzger et al., 2007, 2011, see however Mösta et al. 2014), they necessitate rapid rotation and an extremely strong large-scale poloidal magnetic field $\gtrsim 10^{15}$ G to facilitate jet launching (e.g., Tchekhovskoy & Giannios 2015; Gottlieb et al. 2022a). This raises the question of whether a rapidly rotating star, whose core prior to collapse is predicted to possess a comparatively weak magnetic field $B\approx 10^{10}\,\,{{\rm G}}$ (Aguilera-Dena et al., 2018; Müller, 2024), can generate such strongly magnetized compact objects.
Two possible magnetar formation channels have been proposed (e.g., Spruit 2008; Igoshev et al. 2021 for reviews). In the “fossil field” scenario, an NS acquires its magnetic field externally, from the collapsing star (Ferrario & Wickramasinghe, 2006). If the field has a substantial poloidal component, magnetic flux freezing during the collapse will amplify the field by a factor of $(R_{c}/\,{{R_{\rm NS}}})^{2}$, where $R_{c}$ is the radius of the precollapse star enclosing the NS mass and $\,{{R_{\rm NS}}}$ is the NS radius. This factor can be $\gtrsim 10^{5}$, sufficient to generate an NS magnetic field $\gtrsim 10^{15}\,\,{{\rm G}}$ (see e.g., Makarenko et al., 2021). However, the topology of the core’s magnetic field as predicted by, e.g., the Tayler-Spruit dynamo (TSD; Tayler, 1973; Spruit, 2002; Fuller et al., 2019; Petitdemange et al., 2023a; Skoutnev & Beloborodov, 2024) is primarily toroidal, such that only a small fraction of the total progenitor flux will contribute to the dipole field of the newly formed proto-neutron star (PNS). Furthermore, even if the poloidal field is relatively strong, the small spatial scales that characterize the TSD render the field unlikely to possess a uniform polarity prior to or after the collapse, resulting in a reduced net flux reaching small scales.
@@
At least three mechanisms can be at play over the first few seconds of the PNS cooling evolution: (1) the magnetorotational instability (MRI; Balbus & Hawley, 1991), particularly in the outer layers of the PNS where the angular velocity decreases outward (Akiyama et al., 2003; Obergaulinger et al., 2009; Reboul-Salze et al., 2021); (2) the TSD, particularly in stably stratified regions close to the PNS core where the MRI cannot operate (e.g., Margalit et al. 2022); and (3) an $\alpha-\omega$ dynamo in the convective regions of the PNS (e.g., Duncan & Thompson, 1992; Thompson & Duncan, 1993; Bonanno et al., 2003; Raynaud et al., 2020; White et al., 2022). All three mechanisms are in principle capable of generating fields of strength $\lesssim 10^{16}\,\,{{\rm G}}$ (Thompson & Murray, 2001; Akiyama et al., 2003), though only the MRI and a convective dynamo are likely to dominate near the surface of the PNS (however, see Barrère et al. 2022). The Rossby number $Ro=P/\tau_{c}$ determines the relative contribution of these mechanisms, where $P$ is the PNS spin period and $\tau_{c}$ is the convective turnover time in its envelope. When rotation is slower than the convective motions such that $Ro\gtrsim 1$, the MRI dominates as it grows on the rotation timescale, leading to a stochastic dynamo that generates small-scale multipolar fields (e.g., Guilet et al. 2022; Jacquemin-Ide et al. 2023). On the other hand, for sufficiently rapid rotation $P\lesssim 10$ ms and $Ro\lesssim 0.2$, the MRI is subdominant to the $\alpha-\omega$ dynamo, which can generate a strong ordered dipolar field (Nagakura et al., 2020; Raynaud et al., 2020; White et al., 2022), whose magnitude grows with the rotation rate (Masada et al., 2022).
@@
While the origin of the NS magnetic fields has been extensively debated in the literature, the question of how BHs acquire enough magnetic flux to power relativistic jets in jetted core-collapse events like GRBs has received less attention. Unlike NSs, BHs cannot generate their own fields, which must be inherited externally. One possibility is that the BH acquires its magnetic flux directly through accretion from the infalling progenitor envelope (e.g., Tchekhovskoy & Giannios 2015). However, as we will show, the same arguments raised above against magnetars acquiring large-scale fields from the progenitor star – namely the strength and small coherence scale of the precollapse poloidal field – also apply to the accreting BH case, particularly for stellar cores that retain enough angular momentum to produce GRB jets in the first place. Another possibility is an amplification of the field through dynamo processes in the accretion disk (e.g., Jacquemin-Ide et al., 2023; Most, 2023). This mechanism might also be inefficient for collapsar GRBs given the expected properties of collapsar disks, as we discuss later in light of our stellar evolution models. On the other hand, stellar-mass BHs were previously PNSs, prior to accreting enough mass to undergo gravitational collapse (e.g., Dessart et al. 2012; Aloy & Obergaulinger 2021). This motivates the possibility that collapsar BHs inherit their magnetic fields from their progenitor PNS.
@@
However, this “balding” process may be postponed if the inherited magnetic field can be pinned to the BH by an external current source, such as that supplied by an accretion disk. While not expected to form around an old NS that collapses in isolation (e.g., accretion-induced collapse in a compact binary; Margalit et al. 2015), such a disk may be present around a differentially rotating PNS embedded within the infalling envelope of a rotating star. Nevertheless, it remains unaddressed whether, in astrophysical systems such as collapsars or post-binary NS mergers, the BH can maintain the inherited magnetic flux long enough to launch a sustained relativistic jet necessary to power GRBs.
@@
Here, using an innovative multidisciplinary approach, we explicitly show that massive stars with rapidly rotating cores at the time of collapse generate an accretion disk around the central compact object but do not provide sufficient magnetic flux to power a GRB-like jet. Instead, we propose that such events first create a rapidly spinning PNS surrounded by an accretion disk and that internal dynamo processes amplify the PNS magnetic field to magnetar-level strengths. Then, when the magnetar accretes sufficient mass to collapse, its strong poloidal component is retained by the resulting BH due to confinement by the accretion disk. Therefore, even without ongoing accreted magnetic flux, the BH retains its dipolar field, which can launch the BZ jets that generate GRBs. The astrophysical setup and sequence of events are illustrated in Figure 1.
@@
The paper is organized as follows. In §2, we begin by presenting stellar evolution models that lead to the formation of a PNS with an accretion disk and use them to explain why the intrinsic poloidal flux fed by stellar infall is likely insufficient to power a GRB jet. In §3, we estimate the conditions necessary for the BH to capture the field of the collapsing PNS, which constrain the expected natal spin and magnetic flux acquired by the BH during this process. In §4, we conduct general-relativistic magnetohydrodynamic (GRMHD) simulations demonstrating that BHs retain their magnetic field without further flux accumulation so long as an accretion disk confines them. We conclude in §5, and discuss the implications of our results on GRBs and their associated SNe in §6.
@@
Figure 1: Not-to-scale illustration of the proposed model. Top: a rotating massive star (light blue) collapses, forming a rapidly rotating proto-neutron star (white) surrounded by an accretion disk (dark blue) due to the excess angular momentum of the innermost layers of the stellar core (§2). The magnetic field present in the progenitor star due to the Tayler-Spruit dynamo is too weak and/or small-scale (randomly oriented cyan loops) to contribute a substantial magnetic flux onto the compact object (§2.1). A convective dynamo within the rapidly spinning proto-neutron star efficiently generates a strong, ordered magnetic field (pink lines). Bottom: the proto-neutron star collapses into a moderately spinning black hole (black). The accretion disk surrounding the proto-neutron star at the time of collapse quickly spreads inwards viscously to the black hole’s innermost stable circular orbit, anchoring the proto-neutron star’s magnetic field lines to the BH faster than the “balding” process. This enables the black hole to form with strong magnetic fields already threading through it (§3). The strongly magnetized and moderately spinning black hole then launches a pair of Blandford-Znajek jets (yellow) that break through the outer layers of the collapsing star and power gamma-ray burst emission (§4).
@@ 2 Black hole formation in stellar evolution models
Figure 2: From top to bottom, as a function of the local free fall timescale: specific angular momentum profile (gray band brackets $j_{\mathrm{ISCO}}(a_{\mathrm{BH}}=0)$ and $j_{\mathrm{ISCO}}(a_{\mathrm{BH}}=1)$); mass density (red) and enclosed mass (orange; right vertical axis); estimate of the accretion rate (magenta) and local infall velocity profile (cyan; right vertical axis); B-field components predicted by the Tayler-Spruit dynamo; Alfvén, TSD, and free fall local timescales (purple, yellow, and dashed gray, respectively); and estimated cumulative magnetic flux $\Phi$, where the red band is the required flux to power a GRB jet (see text).
@@
The top panel of Fig. 2 depicts the specific angular momentum of the infalling matter, $j=r\omega^{2}$ (blue), and the innermost stable circular orbit (ISCO) specific angular momentum, $j_{\mathrm{ISCO}}$ (black), as a function of the enclosed mass and angular momentum calculated following Bardeen et al. (1972). In the innermost layers, where $\tau_{\mathrm{free\ fall}}\lesssim 3\,\mathrm{s}$, we find $j\gtrsim j_{\mathrm{ISCO}}$, which prohibits a prompt collapse to a BH and necessitates a phase of a PNS with an accretion disk (see also Dessart et al., 2008). This result is sensitive to uncertain angular momentum transport processes in the star, which are virtually unconstrained in this mass regime. As outlined in Appendix A, we adopt angular momentum transport via TSD in radiative layers (Spruit, 2002), which provides relatively weak transport by magnetic torques (compared to Fuller et al. 2019; Fuller & Lu 2022). In convective layers, angular momentum transport is assumed to be quite efficient, with a diffusion coefficient proportional to the convective diffusivity. We also neglect possible late spin-up of the core due to gravity waves excited by convective shell burning (e.g. Fuller et al., 2015).
@@
Using the same initial conditions and incorporating magnetic torques from [^1] Implemented in https://github.com/MESAHub/mesa-contrib/. Fuller & Lu (2022), the model still exhibits chemically homogeneous evolution. However, most of the angular momentum is lost post-hydrogen core burning, and no layer retains enough angular momentum at the onset of collapse to form an accretion disk. Different assumptions can result in the inner regions spinning less rapidly, allowing a prompt collapse to a BH (MacFadyen & Woosley, 1999). In such cases, one can estimate the BH seed formation time $t_{\mathrm{BH}}$ as the free fall time of the layers with $j<j_{\mathrm{ISCO}}(a(m,j))$, and the accretion disk might form from the outer parts of the stellar envelope. This raises difficulties for this scenario in producing GRB progenitors.
@@
The second panel portrays the mass density profile (red; left vertical axis), and the enclosed mass (orange; right vertical axis). The total mass of the star at the onset of core collapse is $\sim{}25\,M_{\odot}$ due to wind mass loss (see Appendix A). The third panel shows the infall velocity profile (cyan; right axis;=); we evolve the star until it reaches an infall velocity of $-300\,\mathrm{km\ s^{-1}}$. The third panel also shows a rough estimate of the mass accretion rate calculated as $\dot{M}={\rm d}m/{\rm d}\tau_{\mathrm{free\ fall}}$ (magenta). The horizontal blue shaded band highlights $0.1\leq\dot{M}/M_{\odot}\,\,{{\rm s}}^{-1}\leq 3$ (see §3).
@@
The fourth panel depicts the radial ($B_{r}$, pink) and azimuthal ($B_{\varphi}$, green) components of the magnetic field $B$ predicted by the TSD (Spruit, 2002). The region where the field drops below 100 G is convective at the onset of core collapse and thus does not contain TSD-generated fields. These fields are, by definition, local and do not maintain the large-scale structure or a poloidal topology, both of which are required to power electromagnetically driven jets (see §2.1).
@@
$\tau_{\mathrm{TS}}=\tau_{\mathrm{Alfven}}^{2}\times\omega\,,$ (3)
where $\omega$ is the local rotational frequency (e.g. Pitts & Tayler, 1985; Spruit, 1999). This implicitly assumes $\tau_{\mathrm{Alfven}}\gg 2\pi/\omega$. We show the rotational period $2\pi/\omega$ as a dotted gray line in the fifth panel, which illustrates that this assumption holds within the mass that will form the PNS (red vertical dashed line) and in the outermost layers with free-fall times longer than $\simeq 1\,\mathrm{sec}$. However, in between, the assumption of flux freezing does not hold. In the innermost layers, with $m\lesssim M_{\odot}$ and $\tau_{\mathrm{free\ fall}}\lesssim 0.1\,\mathrm{s}$, both $\tau_{\mathrm{Alfven}}$ and $\tau_{TS}$ are significantly longer than the free-fall timescale, justifying the assumption of flux freezing. Moving outward in the region $\,{M_{\odot}}\lesssim m\lesssim 3\,\,{M_{\odot}}$, corresponding to infall times $\tau_{\mathrm{free\ fall}}\lesssim 2\,\mathrm{s}$, the Alfvén and TSD timescales become shorter than the infall timescale. Thus, magnetic diffusion and reconnection could occur, leading to a lower magnetic flux reaching the seed BH (without dynamos).